OpenStudy (anonymous):

show z=f(x^(2)+y^(2)) equals y(dz/dx)-x(dz/dy)=0

OpenStudy (anonymous):

z = f( x^2 + y^2) , use chain rule

OpenStudy (anonymous):

hmmm, is z = f(x,y) ?

OpenStudy (anonymous):

dz/dx = df/ du * du / dx

OpenStudy (anonymous):

where u = x^2 + y^2

OpenStudy (anonymous):

du/dx = 2x + 2y dy / dx

OpenStudy (anonymous):

i needs to equal the second part somehow

OpenStudy (anonymous):

i need to know if this is two variable or one variable calculus

OpenStudy (anonymous):

multivariable / cal 3

OpenStudy (anonymous):

oh , then you have partial derivatives